# Causal Inference with Large Language Model: A Survey

### v2

@@ Abstract
Jing Ma Case Western Reserve University, Cleveland, OH 44106 jing.ma5@case.edu
@@ 1.1 NLP, LLM, and Causality
Figure 1: Representative causal tasks, their positions in the causal ladder, and examples of prompts. PCD = pairwise causal discovery; CA=causal attribution; ATE=average treatment effect; CDE=controlled direct effect; BAJ=backdoor adjustment; CE=causal explanation; CR=counterfactual reasoning; NDE=natural direct effect.
@@ 1.4 Contribution and Uniqueness
Alpaca-7B (Jin et al.), FLAN-T5 (Liu et al.),
BERT-Base, RoBERTa-Base (Gao et al.),
Luminous (Zečević et al.), OPT (Zečević et al.; Liu et al.),
Mistral-7B (Cai et al.), Mistral-7B+Fine-tuning (Cai et al.) , leaf, text width=34em ] ] [ Full Graph Discovery
@@
GPT-3.5 turbo+RAG (Pawlowski et al.) , leaf, text width=34em ] ] [ Causal Effect in Model [ GPT-2, GPT-Neo, GPT-3 Instruct, LLaMA, Alpaca (Stolfo et al.) , leaf, text width=34em ] ] ] [ Other Causal
@@
FLAN-T5 11B (Gao et al.) , leaf, text width=34em] ] ] ]
Figure 2: The major causal tasks and LLMs evaluated for these tasks. Noticeably, the citations in the figure correspond to the work of evaluations, rather than the original work of these models themselves.
@@ 2.2 Causal Tasks and Related Rungs in Ladder of Causation
Causal inference involves various tasks. Figure 1 shows an overview of these tasks and their positions in the ladder of causation. We also show several examples of prompts corresponding to each rung. We list several main causal tasks as follows:
@@
Corr2Cause (2023b) CD 200K (S) Mixed S 1 (2023b)
CLADDER (2023a) Eff, CR, CE 10K (S) Mixed S 1 (2023a; 2023b)
BN Repo (2022) CD 4$\sim$84 (R) Mixed R 8 (2023)
COPA (2011) CD 1,000 (Q) Dailylife R 1 (2023; 2011)
E-CARE (2022) CD, CE 21K (Q) Mixed R 1 (2023; 2022)
@@
CausalNet (2016) CD 62M (R) Mixed S 1 (2016; 2022)
CausalBank (2021) CD 314 M (P) Mixed S 1 (2021; 2022)
WIKIWHY (2022) CD,CE 9K (Q) Mixed R 1 (2022)
@@
CRASS (2022) CR 275 (Q) Mixed R 1 (2022)
CausalQA (2022) CD, CE 1.1M (Q) Mixed R 10 (2024; 2022; 2024)
CALM-Bench (2023) CD, CA 281K (Q) Mixed R 6 (2023)
CausalBench (2024b) Corr, CD 4$\sim$195 (R) Mixed R 15 (2024b)
CaLM (2024a) Rung 1$\sim$3 126K (S) Mixed S 20 (2024a)
@@ 3 Methodologies
Prompting. Most existing works on causal reasoning with LLMs Chen et al. (2024a); Kıcıman et al. (2023); Long et al. (2023); Jin et al. (2023a) focus on prompting, as it is the simplest approach. This includes regular strategies (like basic prompting, In-Context Learning (ICL) Brown et al. (2020), and Chain-of-Thought (CoT) Wei et al. (2022)) and causality-specific strategies Jin et al. (2023a). For regular prompting, basic prompts (i.e., directly describing the question without any example or instruction) are most frequently used. There are also other efforts to devise more advanced prompting strategies. Among them, CaLM Chen et al. (2024a) has tested 9 prompting strategies including basic prompt, adversarial prompt Wallace et al. (2019); Perez and Ribeiro (2022), ICL, 0-shot CoT (e.g., “let’s think step by step” without any examples) Kojima et al. (2022), manual CoT (i.e., guide models with manually designed examples), and explicit function (EF) (i.e., using encouraging language in prompts) Chen et al. (2024a). Other works Kıcıman et al. (2023); Long et al. (2023); Gao et al. (2023); Ban et al. (2023) also design different prompt templates. These works show substantial improvement potential of prompt engineering in causal reasoning tasks. For example, studies Kıcıman et al. (2023); Chen et al. (2024a); Long et al. (2023) highlight that simple phrases like "you are a helpful causal assistant" can significantly boost performance. Additionally, there are causality-specific strategies, such as CausalCoT Jin et al. (2023a), which combine CoT prompting with causal inference principles Pearl and Mackenzie (2018).
@@
Fine-tuning. Fine-tuning, as a widely recognized technique in general LLMs, is now also starting to gain attention for its application in causal tasks. Cai et al. Cai et al. (2023) propose a fine-tuned LLM for the pairwise causal discovery task (PCD, introduced in Section 4.1). This method generates a fine-tuning dataset with a Linear, Non-Gaussian, Acyclic Model Shimizu et al. (2006), uses Mistral-7B-v0.2 Jiang et al. (2023) as LLM backbone, and runs instruction finetuning with LoRA Hu et al. (2021). The results achieve significant improvement compared with the backbone without fine-tuning.
@@
Combining traditional causal methods. A line of studies combines LLMs with traditional causal methods. Considering causal inference often heavily relies on numerical reasoning, an exploration in Ban et al. (2023) leverages LLMs and data-driven causal algorithms such as MINOBSx Li and Beek (2018) and CaMML O’Donnell et al. (2006). This method outperforms both original LLMs and data-driven methods, indicating a promising future for combining the language understanding capability of LLMs and the numerical reasoning skills of data-driven methods in complicated causal tasks. Jiralerspong et al. Jiralerspong et al. (2024) combine LLM with a breadth-first search (BFS) approach for full graph discovery. It considers each causal relation query as a node expansion process, and gradually constructs the causal graph by traversing it with BFS. This method significantly reduces the time complexity from $O(n^{2})$ to $O(n)$, where $n$ is the number of variables. While it does not require access to observational data, their experiments show that the performance can be further enhanced with observational statistics.
Knowledge augmentation. LLMs with augmented knowledge can often better execute tasks for which they are not well-suited to perform by themselves, particularly for causal tasks that require domain-specific knowledge. Pawlowski et al. (2023) introduced two types of knowledge augmentation: context augmentation, which provides causal graphs and ITEs in the prompt, and tool augmentation, offering API access to expert systems for causal reasoning. Tool augmentation performs more robustly across varying problem sizes, as the LLM relies on the API for reasoning instead of reasoning through the graph itself. LACR Zhang et al. (2024) applies retrieval augmented generation (RAG) to enhance the knowledge base of LLM for causal discovery, where the knowledge resources are from a large scientific corpus containing hidden insights about associational/causal relationships. Similar approaches Samarajeewa et al. (2024) employ causal graphs as external sources for causal reasoning. Knowledge augmentation is especially effective for domain-specific tasks, as seen in RC2R Yu et al. (2024) for financial risk analysis and CausalKGPT Zhou et al. (2024a) for aerospace defect analysis.
4 Evaluations of LLMs in Causal Tasks
This section summarizes recent evaluation results of LLMs in causal tasks. We mainly focus on causal discovery and causal effect estimation, and also introduce several representative tasks spanning Rung 1$\sim$3. A collection of datasets used in LLM-related causal tasks is shown in Table 1. In Table 2 and Table 3, we compare the performance of different LLMs in different tasks (including causal discovery and other tasks spanning different rungs) on multiple datasets. The mentioned LLMs include ada, babbage, curie, davinci Brown et al. (2020), text-ada-001, text-babbage-001, text-curie-001, text-davinci-001, text-davinci-002, text-davinci-003 Ouyang et al. (2022), Llama 2 (7B, 13B, 70B) Touvron et al. (2023), and OpenAI’s GPT series Achiam et al. (2023); OpenAI (2022).
4.1 LLM for Causal Discovery
@@
Pairwise causal discovery (PCD). PCD focuses on inferring the causal direction ($A\rightarrow B$ or $A\leftarrow B$) or determining the existence of a causal relationship. Kıcıman et al. (2023) use variable names in prompts, showing that LLMs (e.g., GPT-3.5, GPT-4) outperform state-of-the-art methods on datasets like CauseEffectPairs Mooij et al. (2016) and domain-specific datasets like neuropathic pain Tu et al. (2019). However, some studies Zečević et al. (2023) suggest LLMs often act as "causal parrots", merely repeating embedded causal knowledge. Jin et al. (2023b) proposes a correlation-to-causation inference (Corr2Cause) task, where LLMs performed close to random. Although fine-tuning improved their performance, they still struggle with generalization in out-of-distribution scenarios. In summary, many studies Gao et al. (2023); Kıcıman et al. (2023); Jin et al. (2023b, a); Chen et al. (2024a) take a nuanced stance, acknowledging LLMs’ strengths in PCD tasks while highlighting their limitations in reliably determining the existence of causal relationships.
@@ 4.4 Main Observations and Insights
From the evaluation above and results shown in Table 2 and Table 3, we summarzie the main observations as follows: (1) Model performance: In general, many LLMs exhibit impressive performance in various causal tasks, especially in causal discovery, even with basic prompts. In some cases, their performance can be comparable to or even surpass human-level reasoning Kıcıman et al. (2023). However, as the task difficulty increases from Rung 1$\sim$3, their performance becomes less satisfactory in higher-level complicated causal reasoning tasks Chen et al. (2024a). (2) Enhancement through proper strategies: The performance of LLMs can be significantly enhanced with effective prompting strategies, such as few-shot ICL and CoT. These approaches enable models to leverage contextual information and reasoning processes to achieve better results. Additionally, these models can provide valuable insights through causal explanations. (3) General patterns: While no definitive laws determine model performance universally, certain trends are still observable. For instance, scaling laws suggest that larger models generally perform better, although this is not always that straightforward. These trends provide valuable insights that can guide the future design and development of models. (4) Variability in model effectiveness: There is currently no universally superior LLM or strategy for causal tasks, as their effectiveness can vary significantly depending on the specific scenario. These observations highlight the need for more nuanced and adaptable approaches. (5) Common issues: Current LLMs still struggle with many issues in causal tasks. For example, the answers often lack robustness and are sensitive to changes in prompts Kıcıman et al. (2023); Jin et al. (2023a). Besides, these models frequently default to memorizing and repeating information rather than actual causal reasoning Zečević et al. (2023), which can limit their effectiveness in complex causal scenarios. Besides, LLMs often fail to generate self-consistent causal answers, i.e., the answers from LLMs often conflict with each other. Ongoing debates about whether LLM truly performs causal inference also compel more in-depth analysis.
@@ 6 Limitations
In this survey, it is important to acknowledge certain limitations that shape the scope and focus of our review. First, our analysis is primarily centered on the application of LLMs for causal inference tasks, thereby excluding exploration into how causality is utilized within LLM frameworks themselves. This decision provides a targeted perspective on leveraging LLMs to enhance causal inference methodologies but does not delve into the internal mechanisms or implementations of causal reasoning within these models.

### v3

@@ Abstract
Jing Ma Department of Computer and Data Sciences, Case Western Reserve University jing.ma5@case.edu
@@ 1.1 NLP, LLM, and Causality
Figure 1: Representative causal tasks, their positions in the causal ladder, and examples of prompts w.r.t. mode, question type, and prompting strategy. PCD = pairwise causal discovery; CA=causal attribution; ATE=average treatment effect; CDE=controlled direct effect; BAJ=backdoor adjustment; CE=causal explanation; CR=counterfactual reasoning; NDE=natural direct effect.
@@ 1.4 Contribution and Uniqueness
Alpaca-7B (Jin et al.), FLAN-T5 (Liu et al.), Luminous (Zečević et al.),
BERT-Base, RoBERTa-Base (Gao et al.), OPT (Zečević et al.; Liu et al.),
Mistral-7B (Cai et al.), Mistral-7B+Fine-tuning (Cai et al.),
Causal agent Han et al. (2024), Causal-Copilot Wang et al. (2024) , leaf, text width=34em ] ] [ Full Graph Discovery
@@
GPT-3.5 turbo+RAG, (Pawlowski et al.), Causal agent Han et al. (2024) , leaf, text width=34em ] ] [ Causal Effect in Model [ GPT-2, GPT-Neo, GPT-3 Instruct, LLaMA, Alpaca (Stolfo et al.) , leaf, text width=34em ] ] ] [ Other Causal
@@
FLAN-T5 11B (Gao et al.), Causal-Copilot Wang et al. (2024) , leaf, text width=34em] ] ] ]
Figure 2: The major causal tasks and LLMs evaluated for these tasks. Noticeably, the citations in the figure correspond to the work of evaluations, rather than the original work of these models themselves. The strategies that are not merely based on prompting are highlighted in orange.
@@ 2.2 Causal Tasks and Related Rungs in Ladder of Causation
Causal inference involves various tasks. Figure 1 shows an overview of these tasks and their positions in the ladder of causation. We also show some examples of prompts w.r.t. mode, question type, and prompting strategy. We list several main causal tasks as follows:
@@
Corr2Cause (2023b) CD 200K (S) Math S 1 (2023b)
CLADDER (2023a) Eff, CR, CE 10K (S) Dailylife, etc. S 1 (2023a; 2023b)
BN Repo (2022) CD 4$\sim$84 (R) Health, etc. R 8 (2023)
COPA (2011) CD 1,000 (Q) Commonsense R 1 (2023; 2011)
E-CARE (2022) CD, CE 21K (Q) Commonsense R 1 (2023; 2022)
@@
CausalNet (2016) CD 62M (R) Web text S 1 (2016; 2022)
CausalBank (2021) CD 314 M (P) Web text S 1 (2021; 2022)
WIKIWHY (2022) CD,CE 9K (Q) Wikipedia R 1 (2022)
@@
CRASS (2022) CR 275 (Q) Commonsense R 1 (2022)
CausalQA (2022) CD, CE 1.1M (Q) Web text, etc. R 10 (2024; 2022; 2024)
CALM-Bench (2023) CD, CA 281K (Q) Science, etc. R 6 (2023)
CausalBench (2024b) Corr, CD 4$\sim$195 (R) Health, etc. R 15 (2024b)
CaLM (2024a) Rung 1$\sim$3 126K (S) Commonsense, etc. R & S 20 (2024a)
@@ 3 Methodologies
Prompting. Most existing works on causal reasoning with LLMs Chen et al. (2024a); Kıcıman et al. (2023); Long et al. (2023); Jin et al. (2023a) focus on prompting, as it is the simplest approach. This includes regular strategies (like basic prompting, In-Context Learning (ICL) Brown et al. (2020), and Chain-of-Thought (CoT) Wei et al. (2022)) and causality-specific strategies Jin et al. (2023a). For regular prompting, basic prompts (i.e., directly describing the question without any example or instruction) are most frequently used. There are also other efforts to devise more advanced prompting strategies. Among them, CaLM Chen et al. (2024a) has tested 9 prompting strategies including basic prompt, adversarial prompt Wallace et al. (2019); Perez and Ribeiro (2022), ICL, 0-shot CoT (e.g., "let’s think step by step" without any examples) Kojima et al. (2022), manual CoT (i.e., guide models with manually designed examples), and explicit function (EF) (i.e., using encouraging language in prompts) Chen et al. (2024a). Other works Kıcıman et al. (2023); Long et al. (2023); Gao et al. (2023); Ban et al. (2023) also design different prompt templates. These works show substantial improvement potential of prompt engineering in causal reasoning tasks. For example, studies Kıcıman et al. (2023); Chen et al. (2024a); Long et al. (2023) highlight that simple phrases like "you are a helpful causal assistant" can significantly boost performance. Additionally, there are causality-specific strategies, such as CausalCoT Jin et al. (2023a), which combine CoT prompting with causal inference principles Pearl and Mackenzie (2018). Prompting-based methods can quickly and flexibly adapt to different causal tasks, but are still limited by the specificity of the prompt and thus easily lead to inconsistent causal responses.
@@
Fine-tuning. Fine-tuning, as a widely recognized technique in general LLMs, is now also starting to gain attention for its application in causal tasks. Cai et al. (2023) propose a fine-tuned LLM for the pairwise causal discovery task (PCD, introduced in Section 4.1). This method generates a fine-tuning dataset with a Linear, Non-Gaussian, Acyclic Model Shimizu et al. (2006), uses Mistral-7B-v0.2 Jiang et al. (2023) as LLM backbone, and runs instruction finetuning with LoRA Hu et al. (2021). The results achieve significant improvement compared with the backbone without fine-tuning. However, effective fine-tuning requires large computational resources, and may also suffer from overfitting problems.
@@
Combining traditional causal methods. A line of studies combines LLMs with traditional causal methods. Considering causal inference often heavily relies on numerical reasoning, an exploration in Ban et al. (2023) leverages LLMs and data-driven causal algorithms such as MINOBSx Li and Beek (2018) and CaMML O’Donnell et al. (2006). This method outperforms both original LLMs and data-driven methods, indicating a promising future for combining the language understanding capability of LLMs and the numerical reasoning skills of data-driven methods in complicated causal tasks. Jiralerspong et al. Jiralerspong et al. (2024) combine LLM with a breadth-first search (BFS) approach for full graph discovery. It considers each causal relation query as a node expansion process, and gradually constructs the causal graph by traversing it with BFS. This method significantly reduces the time complexity from $O(n^{2})$ to $O(n)$, where $n$ is the number of variables. While it does not require access to observational data, their experiments show that the performance can be further enhanced with observational statistics. Recently, there have been increasing efforts on causality-driven LLM-based agents such as Causal Agent Han et al. (2024) and Causal-Copilot Wang et al. (2024) that automatically use LLMs to invoke causal tools for causal problems. This line of methods can potentially answer more complex causal questions than traditional methods and plain LLMs, but harmonizing between LLMs and data-driven causal approaches can also be a subtle problem.
@@
Knowledge augmentation. LLMs with augmented knowledge can often better execute tasks for which they are not well-suited to perform by themselves, particularly for causal tasks that require domain-specific knowledge. Pawlowski et al. (2023) introduced two types of knowledge augmentation: context augmentation, which provides causal graphs and ITEs in the prompt, and tool augmentation, offering API access to expert systems for causal reasoning. Tool augmentation performs more robustly across varying problem sizes, as the LLM relies on the API for reasoning instead of reasoning through the graph itself. LACR Zhang et al. (2024) applies retrieval augmented generation (RAG) to enhance the knowledge base of LLM for causal discovery, where the knowledge resources are from a large scientific corpus containing hidden insights about associational/causal relationships. Similar approaches Samarajeewa et al. (2024) employ causal graphs as external sources for causal reasoning. Knowledge augmentation is especially effective for domain-specific tasks, as seen in RC2R Yu et al. (2024) for financial risk analysis and CausalKGPT Zhou et al. (2024a) for aerospace defect analysis. However, high-quality causal knowledge sources are hard to collect, which can limit the LLM performance and also increase the method complexity.
4 Evaluations of LLMs in Causal Tasks
This section summarizes recent evaluation results of LLMs in causal tasks. We mainly focus on causal discovery and causal effect estimation, and also introduce several representative tasks spanning Rung 1$\sim$3. A collection of datasets used in LLM-related causal tasks is shown in Table 1. In Table 2 and Table 4, we compare the performance of different LLMs in different tasks (including causal discovery and other tasks spanning different rungs) on multiple datasets. The mentioned LLMs include ada, babbage, curie, davinci Brown et al. (2020), Qwen (14B) Bai et al. (2023), text-ada-001, text-babbage-001, text-curie-001, text-davinci-001, text-davinci-002, text-davinci-003 Ouyang et al. (2022), Llama 2 (7B, 13B, 70B) Touvron et al. (2023), OpenAI’s GPT series Achiam et al. (2023); OpenAI (2022), Mistral (7B) Jiang et al. (2023), and Claude 2 Models (2023).
4.1 LLM for Causal Discovery
Model Accuracy
GPT-3.5 0.58
GPT-4 0.76
GPT-4-Turbo 0.82
Llama2 (13B) 0.74
Claude 2 0.74
Mistral (7B) v0.2 0.75
Fine-tuned Mistral (7B) v0.2 0.90
Table 3: Accuracy of LLMs and a fine-tuned Mistral in PCD. The results are sourced from Cai et al. (2023).
@@ 4.1 LLM for Causal Discovery
Pairwise causal discovery (PCD). PCD focuses on inferring the causal direction ($A\rightarrow B$ or $A\leftarrow B$) or determining the existence of a causal relationship. Kıcıman et al. (2023) use variable names in prompts, showing that LLMs (e.g., GPT-3.5, GPT-4) outperform state-of-the-art methods on datasets like CauseEffectPairs Mooij et al. (2016) and domain-specific datasets like neuropathic pain Tu et al. (2019). With proper techniques such as fine-tuning (as the example of fined-tuned Mistral shown in Table 3), the performance of PCD can be improved significantly. However, some studies Zečević et al. (2023) suggest LLMs often act as "causal parrots", merely repeating embedded causal knowledge. Jin et al. (2023b) proposes a correlation-to-causation inference (Corr2Cause) task, where LLMs performed close to random. Although fine-tuning improved their performance, they still struggle with generalization in out-of-distribution scenarios. In summary, many studies Gao et al. (2023); Kıcıman et al. (2023); Jin et al. (2023b, a); Chen et al. (2024a) take a nuanced stance, acknowledging LLMs’ strengths in PCD tasks while highlighting their limitations in reliably determining the existence of causal relationships.
@@ 4.4 Main Observations and Insights
From the evaluation above and results shown in Table 2 $\sim$ Table 4, we summarzie the main observations as follows: (1) Model performance: In general, many LLMs exhibit impressive performance in various causal tasks, especially in causal discovery, even with basic prompts. In some cases, their performance can be comparable to or even surpass human-level reasoning Kıcıman et al. (2023). However, as the task difficulty increases from Rung 1$\sim$3, their performance becomes less satisfactory in higher-level complicated causal reasoning tasks Chen et al. (2024a). (2) Enhancement through proper techniques: The performance of LLMs can be significantly enhanced with effective prompting strategies (such as few-shot ICL and CoT) and other techniques like fune-tuning. These approaches enable models to be more causality-focused, with improved ability of leveraging context and adaptively following correct steps in different causal reasoning tasks. Additionally, these models can provide valuable insights through causal explanations. (3) General patterns: While no definitive laws determine model performance universally, certain trends are still observable. For instance, scaling laws suggest that larger models generally perform better, although this is not always that straightforward. These trends provide valuable insights that can guide the future design and development of models. (4) Variability in model effectiveness: There is currently no universally superior LLM or strategy for causal tasks, as their effectiveness can vary significantly depending on the specific scenario. These observations highlight the need for more nuanced and adaptable approaches. (5) Common issues: Current LLMs still struggle with many issues in causal tasks. For example, the answers often lack robustness and are sensitive to changes in prompts Kıcıman et al. (2023); Jin et al. (2023a). Furthermore, these models frequently default to memorizing and repeating information rather than actual causal reasoning Zečević et al. (2023), which can limit their effectiveness in complex causal scenarios. Besides, LLMs often fail to generate self-consistent causal answers, i.e., the answers from LLMs often conflict with each other. Ongoing debates about whether LLM truly performs causal inference also compel more in-depth analysis.
@@ 6 Limitations
In this survey, it is important to acknowledge certain limitations that shape the scope and focus of our review. First, our analysis is primarily centered on the application of LLMs for causal inference tasks, thereby excluding exploration into how causality is utilized within LLMs themselves. This decision provides a targeted perspective on leveraging LLMs to enhance causal inference but does not delve into the internal mechanisms or implementations of causal reasoning within these models.